### v1

# MedMamba: Vision Mamba for Medical Image Classification

Abstract
Medical image classification is a very fundamental and crucial task in the field of computer vision. These years, CNN-based and Transformer-based models are widely used in classifying various medical images. Unfortunately, The limitation of CNNs in long-range modeling capabilities prevent them from effectively extracting fine-grained features in medical images , while Transformers are hampered by their quadratic computational complexity. Recent research has shown that the state space model (SSM) represented by Mamba can efficiently model long-range interactions while maintaining linear computational complexity. Inspired by this, we propose Vision Mamba for medical image classification (MedMamba). More specifically, we introduce a novel Conv-SSM module, which combines the local feature extraction ability of convolutional layers with the ability of SSM to capture long-range dependency. To demonstrate the potential of MedMamba, we conduct extensive experiments using three publicly available medical datasets with different imaging techniques (i.e., Kvasir (endoscopic images), FETAL_PLANES_DB (ultrasound images) and Covid19-Pneumonia-Normal Chest X-Ray (X-ray images)) and two private datasets built by ourselves. Experimental results show that the proposed MedMamba performs well in detecting lesions in various medical images. To the best of our knowledge, this is the first Vision Mamba tailored for medical image classification. The purpose of this work is to establish a new baseline for medical image classification tasks and provide valuable insights for the future development of more efficient and effective SSM-based artificial intelligence algorithms and application systems in the medical. Source code has been available at https://github.com/YubiaoYue/MedMamba.
Keywords Medical Images  $\cdot$ Image Classification $\cdot$ Deep Learning  $\cdot$ State Space Models  $\cdot$ Transformers
1 Introduction
Modern medical research is inseparable from the support of various medical images[1]. Medical images are designed to provide accurate visual representation of the structure and function of various tissues and organs of the human body, to assist medical professionals and scientific researchers in conducting detailed exploration of normal and abnormal conditions in patients’ bodies, thereby serving clinical and research purposes. Purpose. Today, both cutting-edge medical research in laboratories and disease diagnosis by clinicians rely heavily on the rich information provided by medical image analysis to facilitate scientific inference and diagnosis[2, 3].
With the continuous development of medical technology, various medical imaging methods have emerged. In clinical settings, the most widely used medical imaging techniques include computed tomography (CT), ultrasound (US), X-rays, endoscopic imaging, and magnetic resonance imaging (MRI), among others. Unfortunately, the accuracy of detection and diagnosis for cancer and/or many other diseases depends on the professional knowledge of individual clinicians (such as radiologists and pathologists), which leads to significant differences in the way clinicians read and interpret medical images. In order to address and overcome this clinical challenge, many computer-aided detection and diagnosis (CAD) solutions have been developed and tested, aimed at helping clinical doctors more effectively read medical images and make diagnostic decisions in a more accurate and objective manner[4].
Numerous studies and applications have shown that deep learning has great potential in image processing, making it a mainstream technology in the field of CAD and is widely used in medical image classification, medical image segmentation, and medical target detection[5]. It is worth noting that in the clinical diagnosis process, accurately classifying various medical images is the most basic and extremely important task among many tasks[6, 7, 8], because image classification is an important part of patient diagnosis and image labels for classification tasks are easier to obtain. In addition, from the perspective of various algorithms and applications of deep learning, more advanced tasks, such as semantic segmentation, target detection, image generation, and image reconstruction, often directly rely on various data sets (such as the natural image data set ImageNet) Well-designed classification model. These classification models are used to act as backbones, encoders, and decoders in various model architectures for efficient feature extraction from image inputs[9, 10, 11].
A large number of studies have shown that classification models based on CNN and Transformer exhibit excellent performance in various medical image classification tasks and are widely used in various healthcare applications[12, 13, 14, 15, 16]. However, both CNN-based models and Transformer-based models have inherent limitations. CNN-based models are limited by local receptive fields, which greatly hinders their ability to capture long-range information[17, 18]. This often results in insufficient feature extraction, resulting in suboptimal classification results. Although the Transformer-based model shows excellent performance in global modeling, the self-attention mechanism requires quadratic complexity in terms of image size, resulting in a high computational burden, making the model unable to be applied to applications that require dense detection tasks (such as segmentation and reconstruction) and various practical scenarios.
Based on these facts, we are forced to develop a novel medical image classification architecture capable of capturing long-range information effectively and maintaining linear computational complexity. Recently, state space models (SSM) have attracted great interest from researchers. Building on classical SSM research [19], modern SSMs such as Mamba not only establish long-distance dependencies but also exhibit linear complexity related to input size. In addition, SSM-based models have been extensively studied in many fields, including language understanding[20, 21], general vision [18, 22], medical segmentation[23, 24], etc. In particular, some studies have utilized SSM to perform various image processing tasks, such as natural image classification and medical image segmentation. However, to the best of our knowledge, the performance of SSM-based medical image classification models remains to be explored.
Influenced by the success of VMamba [22] in natural image classification tasks, we propose MedMamba for the first time, a simple model based on modern SSM, aiming to demonstrate the potential of Mamba in various medical image classification tasks. Specifically, the core of MedMamba is composed of a module called Conv-SSM. Unlike natural images, there is a high degree of similarity between various medical images, so the model is required to be able to effectively extract both fine-grained features and coarse-grained features. Based on this basic fact, Conv-SSM uses both the local feature extraction ability of the convolutional layer and the long-range modeling ability of SSM to achieve the effective extraction of coarse and fine-grained features in medical images and accurate image classification. Comprehensive experiments were conducted on five medical datasets acquired by different imaging techniques to demonstrate the potential of MedMamba modules in medical image classification. Specifically, we conduct extensive experiments on three public datasets and two private datasets, and the results show that MedMamba can achieve competitive performance.
The main contributions of this article can be summarized as follows: 1) We propose MedMamba, marking the first exploration of the potential application of SSM based models in medical images with different modalities. 2) A comprehensive experiment was conducted on five datasets, and the results showed that MedMamba demonstrated considerable competitiveness. 3) We establish a new baseline for medical image classification tasks and provide valuable insights for the future development of more efficient and effective SSM based artificial intelligence algorithms and application systems in medicine.
2 Methods
2.1 Preliminaries
Among modern SSM-based models, namely the Structured State Space Sequence Model (S4) and Mamba[20, 24, 22], both rely on classical continuous systems that map one-dimensional input functions or sequences, denoted as $x(t)\in\mathcal{R}$, through intermediate implicit states $h(t)\in\mathcal{R}^{N}$ to an output $y(t)\in\mathcal{R}$. The aforementioned process can be represented as a linear Ordinary Differential Equation (ODE):
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t)$ (1)
$y(t)=\mathbf{C}h(t)$
where $\mathbf{A}\in\mathcal{R}^{N\times N}$ represents the state matrix, while $\mathbf{B}\in\mathcal{R}^{N\times 1}$ and $\mathbf{C}\in\mathcal{R}^{N\times 1}$ denote the projection parameters.
S4 and Mamba discretize this continuous system to make it more suitable for deep learning scenarios. Specifically, they introduce a timescale parameter $\mathbf{\Delta}$ and transform $\mathbf{A}$ and $\mathbf{B}$ into discrete parameters $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$ using a fixed discretization rule. Typically, the zero-order hold (ZOH) is employed as the discretization rule and can be defined as follows:
$\overline{\mathbf{A}}=\textup{exp}(\mathbf{\Delta}\mathbf{A})$ (2)
$\overline{\mathbf{B}}=(\mathbf{\Delta}\mathbf{A})^{-1}(\textup{exp}(\mathbf{\Delta}\mathbf{A})-\mathbf{I})\cdot\mathbf{\Delta}\mathbf{B}$
After discretization, SSM-based models can be computed in two ways: linear recurrence or global convolution, defined as equations 3 and 4, respectively.
$h^{\prime}(t)=\overline{\mathbf{A}}h(t)+\overline{\mathbf{B}}x(t)$ (3)
$y(t)=\mathbf{C}h(t)$
$\overline{K}=(\mathbf{C}\overline{\mathbf{B}},\mathbf{C}\overline{\mathbf{AB}},\ldots,\mathbf{C}\overline{\mathbf{A}}^{L-1}\overline{\mathbf{B}})$ (4)
$y=x*\overline{\mathbf{K}}$
where $\overline{\mathbf{K}}\in\mathcal{R}^{L}$ represents a structured convolutional kernel, and $L$ denotes the length of the input sequence $x$.
2.2 MedMamba
As shown in Figure 1, the overall architecture of MedMamba is shown. Specifically, the structure of MedMamba is very simple, including only the patch embedding layer, Conv-SSM Block and patch merging layer.
In the MedMamba, a Patch Embedding layer firstly divides the input image $x\in\mathcal{R}^{H\times W\times 3}$ into non-overlapping patches of size $4\times 4$, thereby mapping the dimensions of the image to $C$ (Here, the default values of $C$ is set to 96), resulting in an embedded image $x^{\prime}\in\mathcal{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$. Subsequently, The $x^{\prime}$ is normalized by using a Layer Normalization [25] before feeding it into the backbone of MedMamba. The backbone is composed of four stages. It should be noted that patch merging layers are used after the output of the first three stages to reduce the height and width of input features, while increasing the number of channels. We use [2, 2, 2, 2] Conv-SSM blocks in four stages, with each stage having [C, 2C, 4C, 8C] channels.
Figure 1: The overall architecture of the MedMamba.
2.3 Conv-SSM Block
Conv-SSM Block is the core module of MedMamba, as shown in Figure 1. Overall, Conv-SSM is a dual-branch module. Firstly, the channel splitting operation is used to split the module input into two sub inputs with the same size. Next, the two sub inputs are fed into two branch (i.e., Conv-branch and SSM-branch)modules, respectively. In the Conv-branch, we simply use continuous convolutional layers to model the local features of the input. In the SSM-branch, we first utilize the layer Normalization layer to process the input. It is worth noting that the input is divided into two branches after passing through Layer Normalization. In the first branch of SSM, the input passes through a linear layer and an activation function. In the second branch of SSM-branch, the input is fed into the 2D Selective Scan module (SS2D) [22] for further feature extraction after passing through a linear layer, depthwise separable convolution, and activation function. Then, Layer Normalization is used to normalize the features and multiply them element by element with the output of the first branch, merging the two branches. Finally, a linear layer is used for feature mixing and added with residual connections to obtain the output of Conv-SSM Block. It is worth noting that the activation function in the SSM-branch defaults to SiLU[26]. At the end of the network, we merge the outputs of two branches along the channel dimension of the feature map and use a convolutional layer with a kernel size of 1 to fuse the feature map, thereby facilitating information interaction between channels.
3 Experiments and results
In this section, we conduct comprehensive experiments on MedMamba for various medical images classfication tasks. Specifically, we evaluate the performance of MedMamba three public medical datasets and two private datasets.
3.1 Datasets
Cervical lymph node lesion ultrasound images (Abbreviated as CLNNUS). CLNNUS is a private dataset containing 3392 cervical lymph node ultrasound images. Specifically, these images were obtained from 480 patients in the Ultrasound Department of the Second Affiliated Hospital of Guangzhou Medical University. The entire dataset is divided into four categories by clinical experts based on pathological biopsy results: normal lymph nodes (referred to as normal), benign lymph nodes (referred to as benign), malignant primary lymph nodes (referred to as primary), and malignant metastatic lymph nodes (referred to as metastatic). The number of normal, benign, primary and metastatic images are 1217, 601, 236 and 1338 respectively.
FETAL_PLANES_DB[27] (Abbreviated as FETAL). A large dataset of routinely acquired maternal-fetal screening ultrasound images collected from two different hospitals by several operators and ultrasound machines. All images were manually labeled by an expert maternal fetal clinician. Images are divided into 6 classes: four of the most widely used fetal anatomical planes (Abdomen, Brain, Femur and Thorax), the mother’s cervix (widely used for prematurity screening) and a general category to include any other less common image plane. Fetal brain images are further categorized into the 3 most common fetal brain planes (Trans-thalamic, Trans-cerebellum, Trans-ventricular) to judge fine grain categorization performance. Based on FETAL’s metadata, we categorize it into six categories. The number of images for each category is as follows: Fetal abdomen (711 images), Fetal brain (3092 images), Fetal femur (1040 images), Fetal thorax (1718 images), Maternal cervis (1626 images), and Other (4213 images) (dataset url: https://zenodo.org/records/3904280).
Covid19-Pneumonia-Normal Chest X-Ray Images[28, 29] (Abbreviated as CPN-CX). Shastri et al collected a large number of publicly available and domain recognized X-ray images from the Internet, resulting in CPN-CX. The CPN-CX dataset is divided into 3 categories, namely COVID, NORMAL and PNEUMONIA. All images are preprocessed and resized to 256x256 in PNG format. It helps the researcher and medical community to detect and classify COVID19 and Pneumonia from Chest X-Ray Images using Deep Learning (dataset url: https://data.mendeley.com/datasets/dvntn9yhd2/1).
Kavsir Dataset[30]. The dataset is collected using endoscopic equipment at Vestre Viken Health Trust (VV) in Norway. The VV consists of 4 hospitals and provides health care to 470,000 people. The Kvasir dataset consists of images, annotated and verified by medical doctors (experienced endoscopists), including several classes showing anatomical landmarks, phatological findings or endoscopic procedures in the GI tract, i.e., hundreds of images for each class. The number of images is sufficient to be used for different tasks, e.g., image retrieval, machine learning, deep learning and transfer learning, etc. The anatomical landmarks include Z-line, pylorus, cecum, etc., while the pathological finding includes esophagitis, polyps, ulcerative colitis, etc. In addition, several set of images related to removal of lesions is also provided, e.g., "dyed and lifted polyp", the "dyed resection margins", etc. The dataset consist of the images with different resolution from 720x576 up to 1920x1072 pixels. Overall, the dataset is divided into the following eight categories: dyed-lifted-polyps (500 images), dyed-resection-margins (500 images), esophagitis (500 images), normal-cecum (500 images), normal-pylorus (500 images), normal-z-line (500 images), polyps (500 images), and ulcerative-colitis (500 images) (dataset url: https://www.kaggle.com/datasets/yasserhessein/the-kvasir-dataset/data).
Large-scale otoscopy dataset [31] (Abbreviated as Otoscopy2024). This dataset is a supplement to previous work. In previous publications, we collected 20542 endoscopic images of ear infections. On this basis, we have added an additional 2039 images from medical institutions. We will name 22581 endoscopic images of the ear as Otoscopy2024. Otoscopy2024 is a large dataset specifically designed for ear disease classification, consisting of 9 categories: Cholestestoma of middle ear(548 images), Chronic suppurative otitis media(4021 images), External auditory cana bleeding (451 images), Impacted cerumen (6058 images), Normal eardrum (4685 images), Otomycosis external (2507 images), Secretory otitis media (2720 images), Tympanic membrane calcification (1152 images), Acute otitis media (439 images).
3.2 Evaluation Metrics
Here, based on the characteristics of medical images, we used accuracy, sensitivity, specificity, precision and F1-score as evaluation metrics for the model. The high or low values of these metrics can reflect the potential of the model. The definitions of these metrics are as follows:
$Overall\ Accuracy=\frac{{TP+TN}}{{TP+TN+FP+FN}}$ (5)
$Precision=\frac{{TP}}{{TP+FP}}$ (6)
$Sensitivity=\frac{{TP}}{{TP+FN}}$ (7)
$Specificity=\frac{{TN}}{{TN+FP}}$ (8)
$F1-score=\frac{{2*Precision*Sensitivity}}{{Precision+Sensitivity}}$ (9)
3.3 Implementation Details
The average value ± standard deviation of the five-fold cross-validation results is used to thoroughly evaluate the potential of MedMamba. Before training the network, we resized all images to 224×224×3. Subsequently, the each image was normalized and standardized. During the training process, we employed the Adam optimizer with a 0.001 initial learning rate, $\beta 1$ of 0.9, $\beta 2$ of 0.999, and weight decay of 1e-4 and Cross-Entropy Loss to optimize the model parameters. To train the network, we utilized the PyTorch framework. We trained the model for 100 epochs and used a batch size of 64. We used an early-stop strategy to prevent model overfitting. Furthermore, we do not apply any data augmentation strategies and pre-trained weights to objectively demonstrate that the results of all model metrics benefit from MedMamba’s unique architecture. The training was conducted on a computer with Ubuntu 22.04 operating system and an NVIDIA GeForce RTX 4090 GPU.
3.4 Main results and findings
In previous work, researchers and developers tended to use classification models carefully designed for ImageNet[32] to classify medical images. Although these models were initially applied to natural image classification, experimental results have shown that they are also effective in medical image classification. Therefore, to demonstrate the potential of MedMamba, we compared it with some state-of-the-art models applied in medical image classification tasks, and the experimental results are shown in Table 1. According to the results, although the proposed MedMamba architecture is very straightforward, it still performs well in medical image classification tasks and surpasses CNNs-based and Transformers-based models in some datasets.
Table 1: The performance of MedMamba on various datasets and comparison with reference models. (Bold font represents the best value. AOA represents the average overall accuracy of the model.)
Dataset Model Fold-1(%)normal-$\uparrow$ Fold-2(%)normal-$\uparrow$ Fold-3(%)normal-$\uparrow$ Fold-4(%)normal-$\uparrow$ Fold-5(%)normal-$\uparrow$ AOA(%)normal-$\uparrow$
CLNNUS ResNet50[33] 80.66 80.75 80.64 80.80 81.43 80.86(±0.29)
ConvNeXt-B [34] 80.78 81.68 81.61 80.64 81.10 81.16(±0.42)
ViT-B [17] 82.57 79.76 78.73 76.96 77.63 79.13(±1.97)
Swin-T [35] 80.16 79.62 80.93 80.10 80.58 80.28(±0.44)
MedMamba 85.38 83.90 83.60 84.49 84.94 84.46(±0.65)
Fetal ResNet50[33] 89.87 89.27 89.10 88.70 89.23 89.23(±0.38)
ConvNeXt-B [34] 89.02 88.66 88.62 88.58 89.63 88.89(±0.40)
ViT-B [17] 90.39 90.15 89.10 91.81 88.22 89.93(±1.22)
Swin-T [35] 90.56 90.64 89.87 90.44 89.47 90.20(±0.45)
MedMamba 93.10 92.82 92.37 93.18 93.77 93.05(±0.46)
CPN-CX MobileNetV3-L[36] 93.62 93.71 93.02 93.70 94.33 93.67(±0.42)
EfficientNetV2-M [37] 95.40 94.91 95.55 95.30 95.57 95.35(±0.24)
ViT-B [17] 94.29 94.35 94.09 94.76 94.41 94.38(±0.22)
Swin-T [35] 95.60 95.89 96.08 95.60 95.99 95.83(±0.20)
MedMamba 96.46 97.13 96.27 96.75 96.67 96.66(±0.29)
Kavsir ResNet50[33] 73.91 74.62 74.94 73.94 74.79 74.44(±0.43)
ConvNeXt-B [34] 74.63 74.23 74.02 73.99 73.59 74.09(±0.34)
ViT-B [17] 73.86 73.99 73.60 73.65 74.00 73.82(±0.17)
Swin-T [35] 74.71 75.04 75.75 74.66 74.49 74.93(±0.45)
MedMamba 74.25 74.63 74.63 73.38 74.50 74.28(±0.47)
Otoscopy2024 ResNet50 [33] 87.57 88.74 88.59 86.62 87.04 87.71(±0.84)
ConvNeXt-B [34] 90.94 89.64 87.88 88.09 88.15 88.94(±1.18)
Vit-B [17] 89.37 92.14 91.13 91.77 90.28 90.94(±1.00)
Swin-T [35] 90.10 89.66 88.87 89.90 90.52 89.81(±0.55)
MedMamba 91.36 90.55 90.05 89.60 90.52 90.41(±0.58)
4 Conclusion
In this work, we first introduce an SSM-based model into medical image classification tasks and propose MedMamba as a baseline. In order to enable MedMamba to effectively model various medical images with fine-grained features, we constructed MedMamba using a hybrid module with convolutional layers and SSM. In order to fully demonstrate that MedMamba’s potential in medical classification tasks directly benefits from SSM, we did not use any pre-training strategies or data augmentation strategies. Extensive experiments were conducted on MedMamba using a large number of medical datasets with different modalities. The results show that MedMamba has excellent performance in medical image classification tasks. In addition, compared to various architectures widely used in medical classification tasks, MedMamba demonstrates strong competitiveness.
In addition, we summarize our future work into the following points: 1) We will further explore and test the potential of MedMamba on medical datasets obtained from other imaging technologies, and optimize the internal architecture of MedMamba. 2) We will further use explainable artificial intelligence to analyze MedMamba’s decision-making mechanism. In addition, the impact of MedMamba’s inference speed and model parameter size on practical applications should also be studied. 3) Due to the advantages of SSM in efficient long sequence modeling, further research on the performance of MedMamba in some high-resolution medical images (such as pathological images) may be beneficial. 4) We will attempt to use the backbone of MedMamba as an encoder or decoder to explore its application potential in more advanced medical image tasks, such as medical image segmentation, medical object detection, medical image registration, and medical image reconstruction.